find the exact value of each function.
1. Cos 7pi/4
2. Sin (-30')
Find the reference angle for the angle with the given measure
1. 13pi/8
2. -7pi/4

Answer by stanbon(75887)

find the exact value of each function.
1. Cos 7pi/4 = cos (-1/4 pi) = cos (1/4 pi) = (sqrt(2))/2
-------------------------
2. Sin (-30') = -sin30 = -1/2
-------------------------
Find the reference angle for the angle with the given measure
1. 13pi/8: ref. angle= (3/8)pi
--------------
2. -7pi/4; ref. angle = (1/4)pi
========================
Cheers,
Stan H.